Gucchi:

How many solutions does the equation 4n + 6 = 4 + 2 + 4n have? Two Infinitely many None One

Hero:

@Gucchi what happens if you subtract \(4n\) from both sides?

Gucchi:

What do you mean?

Gucchi:

I can't figure it out

Hero:

Have you tried subtracting \(4n\) from both sides?

Gucchi:

I did try. I can't seem to get the answer tho

Hero:

What was the result when you subtracted 4n from both sides? Mind posting it here?

Gucchi:

I can't post it because i am on a desktop right now and it doesnt have a cam. I couldn't figure it out when i tried doing it. I'm very confused.

Shadow:

Gucchi are you still taking the exam?

Gucchi:

No, this is my homework. I finished the exam. I got a score of 95%

Shadow:

Nice, good job

Hero:

Exam? What?

Gucchi:

I was taking an exam a while back.

Gucchi:

Hmm

Hero:



Hero:

Can @Gucchi finish that?

Gucchi:

Let me try

Gucchi:

6n on both sides when subtracting by 4n?

Hero:

If you subtract 4n on both sides, how do you end up with 6n on both sides? Can you show your work?

Gucchi:

I added 4n + 6 and i got 10n. I then added 4n + 2 + 4 and i got 10n. I subtracted 10n - 4n on both sides and i got 6n. I don't know if i did it right.

Gucchi:

@Hero

Hero:

You can only add two numbers together if they are like terms (meaning they have the same variable). Does 4n and 6 have the same variable?

Gucchi:

No, it doesn't

Hero:

Exactly.

Hero:

It really doesn't matter in what order you try to solve the problem, you should end up with the same number on both sides. What does that imply?

Gucchi:

I'm not sure. I'm feel so dumb ):

Hero:

Let's start small. What is 4 + 2?

Gucchi:

6

Hero:

Okay so performing that operation yields 4n + 6 = 6 + 4n

Gucchi:

Okay

Hero:

Can you draw any conclusions from there?

Hero:

Let's do this: What happens if n = 1?

Gucchi:

it would be 4 x 1?

Hero:

Okay and what is the resulting equation?

Gucchi:

4n?

Hero:

If n = 1, then 4(1) + 6 = 6 + 4(1)

Gucchi:

Oh yh yh

Hero:

Can you continue to simplify it from there?

Gucchi:

Okay

Gucchi:

4 + 6 = 6 + 4

Gucchi:

Do you add the "n"

Hero:

Nope. you just add the 4 and 6

Hero:

And what is the resulting equation?

Gucchi:

10 = 10

Hero:

Very good. Now what happens if n = 2?

Gucchi:

Then it would be 10(2) = 10(2)?

Hero:

Correct. And what is the resulting equation?

Gucchi:

20 = 20

Hero:

Very good. Now let n = any random number. What is the result?

Gucchi:

20(5) = 20(5). 100 = 100

Hero:

Again you get the same number on both sides. What happens if n = 1000? Do we get the same thing on both sides?

Gucchi:

Yes, you do.

Hero:

So which option is correct?

Gucchi:

100(1000) = 100(1000) 100,000 = 100,000

Hero:

so which option is correct?

Gucchi:

For my question right?

Gucchi:

2nd option

Gucchi:

It makes the most sense

Hero:

Correct finally.

Gucchi:

I finally get it

Gucchi:

Thanks so much!